OpenStudy (anonymous):

2x2 - 12xy - 32y2 = 2(x - 8y)(x + By) B= ____________

OpenStudy (anonymous):

Factor out a 2.

OpenStudy (anonymous):

im pretty sure its 4 but im not sure

OpenStudy (anonymous):

2(x^2 - 6xy - 16y^2)

OpenStudy (anonymous):

its not multiple choice

OpenStudy (anonymous):

And 4 is incorrect..factor \(\sf x^2 - 6xy - 16y^2\), can you do that?

OpenStudy (anonymous):

yes it could with the -8 it makes sense

OpenStudy (anonymous):

anchor man. Is it REALLY you?!

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what now?

OpenStudy (anonymous):

They already give you (x - 8y)..finish factoring \(\sf x^2 - 6xy - 16y^2\)

OpenStudy (anonymous):

his picture is the guy off of the movie anchor man

OpenStudy (anonymous):

by odens beard

OpenStudy (anonymous):

where did you get the 6 from

OpenStudy (anonymous):

We factored out a 2. 2x^2 / 2 = x^2 -12xy / 2 = -6xy -32y^2 = -16y^2

OpenStudy (anonymous):

-32y^2 / 2*

OpenStudy (anonymous):

so it would be 16

OpenStudy (anonymous):

16x^2

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (anonymous):

-8 * ? = -16

OpenStudy (anonymous):

That will be your answer^

OpenStudy (anonymous):

your confusing me now so its -8

OpenStudy (anonymous):

-8 * -8 = -16?

OpenStudy (anonymous):

no its 64 duh

OpenStudy (anonymous):

Yes, so -8 is incorrect. Multiply what to -8 gives us -16?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Yes, so that's our answer.

OpenStudy (anonymous):

ohh ok thanks appreciate it